OpenStudy (anonymous):

CALC II I have a question. The question asks to solve for k for the equation: 3^(k/14)=a We are using logarithms and exponentials but I just don't understand how to isolate k using that? please help!

OpenStudy (amistre64):

logs undo exponents

OpenStudy (amistre64):

log both sides \[ln(3^{k/14})=ln(a)\] exponents pull out of logs \[(k/14)ln(3)=ln(a)\] divide off the ln(3) and multiply thru my 14

OpenStudy (anonymous):

i know so i use log or natural log (ln)? like do I use e^ln(k/14)??? like i dont get it for some reason

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

I SEEE

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

cool

OpenStudy (anonymous):

wow that was extremely stupid. I should have figured that out. thank you again! If i have any more questions I'll let you know!

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

actually sorry I did type this into my webwork problem earlier and it said it was incorrect

Parth (parthkohli):

Just a little interruption, but isn't this Pre-Calculus? :|

OpenStudy (anonymous):

I put in 14[(ln(a))/(ln(3))] and it said incorrect

OpenStudy (anonymous):

its a preview of exponential functions we are learning in calc II. so we know how to derive and integrate. anyway i dont get it

Parth (parthkohli):

Another way to do it is,\[a^b = c \implies \log_a c = b\]\[3^{\large {k\over 14} }=a \implies \log_{3}a = {k \over 14} \implies 14\log_3 a = k\]

OpenStudy (anonymous):

okayy i'll try that out! thank you

Parth (parthkohli):

Well, the above still leads to the answer that @amistre64 provided, but still... :)

OpenStudy (anonymous):

did I do the other way right though? it seems like it was correct. Im not sure why it says incorrect

OpenStudy (amistre64):

you didnt simplify the results

OpenStudy (amistre64):

\[k=14\frac{ln(a)}{ln(3)}\] by change of base rule we get \[k=14~log_3(a)\]

OpenStudy (anonymous):

I also tried that answer to simplify it "14 log3(a)" and it still came out as incorrect Im totally lost

OpenStudy (amistre64):

are you sure you have the question correct?

OpenStudy (anonymous):

yess it says solve for k in the equation 3exp (k/14)=a

OpenStudy (amistre64):

\[3^{\frac{14log_3(a)}{14}}=a\] \[3^{log_3(a)}=a\] \[a=a\]

OpenStudy (amistre64):

umm, exp(n) has a specific meaning; its means e^(n), eulers number

OpenStudy (anonymous):

so I did it wrong? I just thought it meant to the power of

OpenStudy (amistre64):

lets try it this way:\[3e^{k/14}=a\] \[e^{k/14}=a/3\] \[ln(e^{k/14})=ln(a/3)\] \[(k/14)~ln(e)=ln(a/3)\] \[k/14=ln(a/3)\] \[k=14~ln(a/3)\]

OpenStudy (amistre64):

ln(a/3) can also be written as: ln(a) - ln(3) but i cant be sure which one is the "simplified" version of it

OpenStudy (anonymous):

yes it's correct! thankyou so much. Now that I look at it, it's relatively simply lol. wow I feel so stupid but thank you!

OpenStudy (amistre64):

being able to read the notation is useful ;) good luck

OpenStudy (anonymous):

yeah If it had just said e^(whatever) I would have spent a lot less time on that lol